OpenStudy (mahria):

Can somebody please help me out with these 4 math problems? Find the zeros of each quadratic function a.) f(x)=x^2-16 b.)f(x)=x^2-7x+10 c.)f(x)=x^2-5x-24 d.) f(x)=2x^2-7x-4 (^2 means squared)

OpenStudy (danjs):

you can factor to linear terms to make it easy to solve when x=0 or just use the quadratic formula for x , plug in method for quadratic y=a*x^2+b*x+c = 0

OpenStudy (mathmate):

There are different ways to go about solving these equations: Method 1: Factorize. Use these standard identities, f(x)=x^2-a^2=(x+a)(x-a) which means that the zeroes are x=a or x=-a. Example: f9x)=x^2-25=(x+5)(x-5) => x=5 or x=-5 There is a very detailed tutorial on factoring prepared by @bloomlocke367 here: http://openstudy.com/study#/updates/55acffcee4b071e6530c96ce Method 2: use the quadratic formula for the standard equation f(x)=ax^2+bx+c \(\LARGE x=\frac{-b\pm\sqrt{b^2-4a}}{2a}\) Zeroes of all of the above functions can be obtained using the formula.

OpenStudy (mathmate):

* replace 4a by 4ac

OpenStudy (mahria):

so if i use method 2, would the answer i get be the zeros?

OpenStudy (mathmate):

Yes. For example, to solve f(x)=x^2-11x+28 we have a=1, b=-11, c=28 substitute in formula: \(\LARGE x=\frac{11\pm\sqrt{11^2-4(1)(28}}{2(1)}\) \(\LARGE =\frac{11\pm\sqrt{121-112}}{2}\) \(\LARGE =\frac{11\pm 3}{2}\) \(=7~ or~ 4\) so x=7 or x=4

OpenStudy (mahria):

Ohhh

OpenStudy (mahria):

Thank You

OpenStudy (mathmate):

You're welcome!

OpenStudy (bloomlocke367):

Haha thanks @mathmate for the mention :D